OpenStudy (anonymous):

For which equation would you subtract 3, and then multiply by 9 to solve? A.nine times k plus three equals forty B. the quantity k over nine all plus three equals forty C.nine times k minus three equals forty D.the quantity k over nine all minus three equals forty

OpenStudy (anonymous):

B. The equation of b is k/9+3=40 According to the question, subtract both sides by 3:- k/9=37 Multiply it by 9:- k=37*9=333

OpenStudy (anonymous):

A becomes 9k+3=40 c becomes 9k-3=40 D becomes k/9-3=40

OpenStudy (anonymous):

thats what i really need help with

OpenStudy (anonymous):